I'm having trouble solving this equation: 3=7(4-2u)-6u
this is what i have so far
3=7(4-2u)-6u
3=28-14u-6u
then i get lost i really dont know how to solve these equations, please help! thank you

Answer by rwm(914)

continue
combine like terms
3=28-20u
add 20 u and subtract 3
20u=25
u=25/20
u=5/4
u=1.25 or 1 1/4